A new landowner has a triangular piece of flat land she wishes to fence. Starting at the first corner, she measures the first side to be 4.2 m long and is directed 1 radians north of east. From the second corner, the second side is 5.5 m long and is directed 0.2 radians west of north. What is the length of the third side of the fence?
Answer #1

A = 4.2 ( cos 1 i + sin 1 j) = 2.27 i + 3.534 j

B = 5.5 ( - sin 0.2 i + cos 0.2j) = - 1.1 i + 5.39 j

now in roder to have final displacement

A + B + C = 0

(2.27 - 1.1) i + ( 3.534 + 5.39) j + C= 0

C = - 1.6i - 8.92j

magnitide

C = 9.06 m

=======

Comment in case any doubt, will reply for sure.. Goodluck